N2+3h2=2nh3 I mol of nitrogen and 4 mol of hydrogen are taken in 15l flask at27°c after complete conversion of n2to nh3 5 l of h 2o is added pressure in flask?

Grade:11

3 Answers

Vikas TU
7 years ago
N2+3h2=2nh3
Here N2 is the limiting Reagent.
Therefore, moles of NH3 is=> 1*2 => 2 moles.
Now H2O is added,
Hence the reacn. becomes,
NH3  +   H2O  --------------->        NH4OH
2 mol                                           2 mol.
Hence Total pressure in flask woould be,
P*15 = 2*R*300
P = 600R/15 => 40R atm
      or
 => 40 * 0.0821
 => 3.284 atm.
Aditya Singh
4 years ago
Here, 2 moles of NH3 is formed and 1 mole of H2 is leftover.
Also, since 5 litre of water is added, Vol=10L.
P= (1*0.0821*300)/10=2.463atm
penuelzechariah
3 years ago
N2+3h2=2nh3
Here N2 is the limiting Reagent. 
5lit of water is added so volume gets used ,so available space is 15 -5 =10 lit   
ammonia (gas) which has presssure ,but when water is added it gets converted to nh4oh solution (liq)  
 
NH3  +   H2O  --------------->        NH4OH  
 
initial               1mol(LR)           4mol          0mol                            N2                   +   3h2       =   2nh3
final                 0 mol                   8/3mol       2 mol   
 
 
total no of moles after complete conversion is 8/3  nh3 is used up by water   
preeesure in flask is =nrt/v 
                                  =8/3*0.0821*300//  10 
                                  =3.284atm     
total no of moles